OpenStudy (anonymous):

What is the simplified form of (–3x3y2)(5xy–1)? a.)–15x3y2 b.)–15x4y c.)15x4y

OpenStudy (anonymous):

you mean \[-3x ^{3}y ^{2}(5xy -1)\]

OpenStudy (anonymous):

Yess. I forgot to put the "^" mark to show that. Thanks. lol.

OpenStudy (anonymous):

And same for the answer choices. lol...

jimthompson5910 (jim_thompson5910):

-3 times 5 = -15 so it's either A or B to multiply the variable terms, add the corresponding exponents x^3 times x = x^3 times x^1 = x^(3+1) = x^? y^2 times y^(-1) = y^(2 + (-1)) = y^(2 - 1) = y^?

OpenStudy (anonymous):

Thankyou.